OpenStudy (anonymous):

Prove the following: If a is divisible by 3 and b is divisible by 4 then a × b is divisible by 6.

zepdrix (zepdrix):

Hey :)

zepdrix (zepdrix):

So if a is divisible by 3, it means that a is equal to some integer multiple of 3s. \(\large\rm a=3k,\qquad k\in\mathbb Z\)

zepdrix (zepdrix):

We can do similarly with b, it's some amount of 4s, \(\large\rm b=4\ell,\qquad \ell\in\mathbb Z\)

zepdrix (zepdrix):

So then our product,\[\large\rm ab\quad=\quad (3k)(4\ell)\quad=\quad 12k\ell\]Hmm do you see the next step? :)

OpenStudy (anonymous):

yes, thank you! 12kl is divisible by 6

zepdrix (zepdrix):

Good! \(\large\rm =6(2k\ell)\) Since k and l are integers, 2kl is an integer as well, so it works out! Yay team! ୧ʕ•̀ᴥ•́ʔ୨

OpenStudy (anonymous):

thank you very much :)